# Unlocking Personalized Knowledge in Federated Large Language Model: The Power of Mixture of Experts

### v1

@@ Appendix A Algorithm Framework
We now present a detailed overview of our FLEx (Federated LLMs with Personalized Experts) algorithm framework. As illustrated in Algorithm 1, we incorporate personalized expert selection, execute local model updates for each client, and aggregate the attention-based updates at the server level. This process effectively balances global aggregation and local personalization. Algorithm 1 outlines the FLEx framework, beginning with personalized expert selection for each client. Each client utilizes their respective local dataset to perform expert selection, subsequently updating their local model using a personalized MoE layer. During global aggregation phases, clients transmit only their attention layer LoRA adapters to the central server, retaining personalized MoE layers locally. These adapters are aggregated across all clients, and the resulting global attention layer is integrated locally within each personalized MoE model.
@@
8:         Send attention layer LoRA adapters to server
9:         Keep personalized MoE layer locally
@@
11:     Aggregate global attention LoRA adapters
12:     Send global attention to clients and integrate into personalized models locally
@@
Algorithm 2 provides the detailed procedure for personalized expert selection. Specifically, we employ a reconstruction loss strategy to prune the original MoE layers down to a tailored subset of experts uniquely optimized for each client. The selected subset effectively minimizes the discrepancy between the outputs of the full expert set and those produced by the personalized subset, thus ensuring optimal local performance.
@@
3:for each MoE layer independently do
@@
7:     Select personalized expert subset $S_{i}$ by solving:
$S_{i}=\arg\min_{S\subset\{1,\dots,K\},|S|=n}\left(\frac{1}{|\mathcal{D}_{i}|}\sum_{x\in\mathcal{D}_{i}}\|F_{S}(x)-F(x)\|_{F}\right)$
8:     Prune the MoE layer by retaining only experts in $S_{i}$
@@ Appendix B Discussion
Our method retains the pre-trained weights in the MoE layer while adding a personalized side expert. This design maximally preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.
@@
Our method adopts a layer-wise greedy pruning strategy, which ensures minimal reconstruction loss at each individual layer. However, this localized optimization does not necessarily guarantee a globally optimal pruning configuration across the entire model. As a result, the pruning decisions made at early layers may limit the potential of subsequent layers, leading to suboptimal global performance. Exploring more globally-aware pruning strategies across layers that better balance efficiency and model fidelity remains an important and underexplored direction for future research.
@@
It is worth mentioning that [15] proposed a federated learning approach specifically targeting frameworks utilizing Switch Transformers to address data heterogeneity. However, their approach involves fine-tuning all MoE layers, inevitably leading to significant forgetting from the pretrained models. This forgetting issue is particularly concerning given that current decoder-only large language models are pretrained on billions of tokens, making the preservation of pretrained knowledge an important task in federated learning. In contrast, our proposed method maintains the integrity of the original pretrained model parameters by introducing a specialized side-expert specifically designed to handle data heterogeneity. Our approach effectively mitigates the forgetting issue inherent in methods that fine-tune all MoE layers. Furthermore, we validate the applicability and efficacy of our federated learning approach through extensive experiments on billion-token pretrained, instruction-tuned large language models, including Qwen and DeepSeek.
@@ C.1 Evaluation Across Different Model Architectures
FLEx (Ours) 50.33 20.11 21.29 38.93 32.66
C.2 Experiments on Trainable Modules
To further validate the effectiveness of FLEx’s training and sharing strategies, we conducted additional experiments investigating various combinations of trainable modules. Specifically, we explored the following approaches:
  1. 1.
    We select a single expert randomly at each layer for keeping same trainable parameters with FLEx.
  2. 2.
    We select all experts at each layer instead of choosing a single pruned expert, which consequently leads to higher communication costs and an increased number of trainable parameters.
The experimental results, summarized in Table 6, demonstrate that our proposed method achieves comparable performance while significantly reducing both communication and computational costs.
Table 6: Rouge-L performance comparison of different federated learning strategies, including our proposed method, on the Databricks-dolly-15k dataset with Qwen1.5-MoE-A2.7B under a pathological non-IID scenario. The best result over federated method is bolded and the second-best result is underlined. CP: Communication Parameter, TP: Trainable Parameter. denote that further shared and averaging all LoRA layers on MoE layers, which lead to more communication costs during training.
Databricks-dolly-15k
Method Modules CLF CQA IE Summ Avg CP(%$\downarrow$ TP(%$\downarrow$
MoE + FedAvg Single Expert 50.83 34.34 36.26 41.59 40.75 0.1327 0.1327
MoE + FedAvg All Experts 51.80 35.91 38.91 42.12 42.18 3.4172 3.4172
MoE + FedAvg Single Expert 51.48 34.77 37.23 41.74 41.30 0.0672 0.1327
MoE + FedAvg All Experts 51.10 32.51 41.39 40.60 41.40 0.0672 3.4172
FLEx (Ours) Single Experts 52.10 38.23 40.13 42.09 43.13 0.0672 0.1307
C.3 Comparison with Traditional Federated Learning
@@
C.4 Experiments on Expert Aggregation Strategies
In this section, we evaluate several expert aggregation strategies within FLEx. In this paper, FLEx employs a sigmoid function as presented in formula (8). We have experimented with various aggregation strategies for side experts, comparing the sigmoid function against alternative approaches, such as ReLU and Tanh functions. Specifically, the ReLU-based weighting for side expert $g_{e}^{t}$ can be expressed as:
$g_{e}^{t}$ $=\text{ReLU}(\text{router}_{e}(u_{t}^{l})).$ (9)
$g_{e}^{t}$ $=\text{Tanh}(\text{router}_{e}(u_{t}^{l})).$ (10)
As shown in Table 7, our sigmoid-based aggregation method consistently demonstrates superior effectiveness. In contrast, the use of ReLU and Tanh functions results in suboptimal performance across aggregation strategies. From this result, we discover that the aggregation strategy of side expert requires carefully chosen.
Databricks-dolly-15k
Method CLF CQA IE Summ Avg
ReLU 51.16 33.02 37.51 41.78 40.86
Tanh 41.40 20.40 31.34 37.88 32.75
Sigmoid 52.10 38.23 40.13 42.09 43.13
@@ C.4 Experiments on Expert Aggregation Strategies
C.5 Experiments on Expert Load Balancing
We conducted experiments to evaluate the activation frequency of experts within the Qwen1.5-MoE-A2.7B model using the C4 dataset [^4] https://huggingface.co/datasets/allenai/c4 across different methods. Figure 4 illustrates the activation count for individual experts. This analysis demonstrates that our method achieves a more balanced workload distribution across different experts. Specifically, the original model and the FedAvg method tend to produce uneven workloads, causing certain experts, such as expert 31 and expert 43, to become significantly overloaded. Specifically, the original model exhibits a standard deviation of 1859.80, while FedAvg demonstrates a slightly higher imbalance with a standard deviation of 1906.85. In contrast, our proposed FLEx method significantly improves load balancing among experts, achieving a notably lower standard deviation of 1259.88.
C.6 Data Distribution for Non-IID
@@
Figure 5: Distribution of instruction types across different clients in federated learning under various settings. The first panel shows the pathological distribution, where each client has a highly imbalanced data distribution. The second panel illustrates the Dirichlet distribution with alpha=0.1, resulting in a more skewed distribution across clients. The third panel presents the Dirichlet distribution with alpha=1.0, demonstrating a more uniform distribution of categories (classification, closed question answering, information extraction, and summarization) across clients.
C.7 Overview of Hyperparameters
@@
C.8 Dataset Information

### v2

@@ Appendix A Algorithm Framework
We now present a detailed overview of our FLEx (Federated LLMs with Personalized Experts) framework. The cornerstone of FLEx lies in this separation: the personalized experts and routers remain strictly local, preserving client personalization. In contrast, the non-expert layers are aggregated globally, enabling shared knowledge to propagate across the federation. This design allows FLEx to effectively balance powerful personalization with collaborative learning. The process begins with a one-time personalized expert selection for each client (detailed in Algorithm 2). Subsequently, the training proceeds in iterative rounds, as outlined in Algorithm 1. Each round consists of two primary steps:
  1. 1.
    Local Training: Each client finetunes its model by training all non-expert layers and its personalized subset of experts within the MoE layers. The client-specific router is also updated during this step.
  2. 2.
    Global Aggregation: After local training, clients transmit only the parameters of the non-expert layers to the server. The server performs federated averaging (FedAvg) on these parameters and sends the aggregated global non-expert layers back to the clients, who integrate them into their local models.
@@
8:   Send parameters of non-expert layers to server
9:   Keep personalized MoE layer and routers locally
@@
11:  Aggregate parameters of non-expert layers
12:  Send aggregated non-expert layers to clients and integrate into personalized models locally
@@
Algorithm 2 provides the detailed procedure for the initial personalized expert selection. Specifically, we employ a reconstruction loss strategy to prune the original MoE layers down to a tailored subset of experts uniquely optimized for each client. The selected subset effectively minimizes the discrepancy between the outputs of the full expert set and those produced by the personalized subset, thus ensuring strong local performance.
@@
3:for each MoE layer $l$ independently do
@@
7:  Select personalized expert subset $e_{i,l}^{*}$ by solving:
$e_{i,l}^{*}=\arg\min_{e\in\{1,\dots,E\}}\sum_{\mathbf{u}_{j}^{l}\in\mathcal{D}_{i}}\left\|\text{FFN}_{e}(\mathbf{u}_{j}^{l})-\sum_{k=1}^{E}g_{k,j}\cdot\text{FFN}_{k}(\mathbf{u}_{j}^{l})\right\|_{F}^{2},$
8:  Prune the MoE layer by retaining only experts in $e_{i,l}^{*}$
@@ Appendix B Discussion
Our method retains the pre-trained weights in the MoE layer while adding a personalized side expert. This design effectively preserves pre-trained knowledge, leverages personalized experts to enhance client-specific capabilities, benefits from shared global knowledge aggregation, and significantly reduces communication overhead.
@@
A key limitation of FLEx is the restriction to grafting only a single personalized expert for each client per layer. As discussed in Appendix C.2, while a straightforward extension would be to greedily select multiple experts, this approach would introduce substantial computational costs without guaranteeing an optimal combination. Therefore, developing more effective pruning algorithms that can efficiently identify a globally optimal set of personalized experts is a direction for future research.
@@
It is worth mentioning that Mei et al. (2024) proposed a federated learning approach specifically targeting frameworks utilizing Switch Transformers to address data heterogeneity. However, their approach involves fine-tuning all MoE layers, leading to significant forgetting from the pretrained models. This forgetting issue is particularly concerning given that current decoder-only large language models are pretrained on billions of tokens, making the preservation of pretrained knowledge an important task in federated learning. In contrast, our proposed method maintains the integrity of the original pretrained model parameters by introducing a specialized side-expert specifically designed to handle data heterogeneity. Our approach effectively mitigates the forgetting issue inherent in methods that fine-tune all MoE layers. Furthermore, we validate the applicability and efficacy of our federated learning approach through extensive experiments on billion-token pretrained, instruction-tuned large language models, including Qwen and DeepSeek.
Efficiency Analysis.
Although our method introduces an additional personalized expert, it significantly reduces training overhead compared to conventional fine-tuning by restricting backpropagation exclusively to this new expert and its adaptive gate. Moreover, the impact on inference latency is negligible, as the grafted expert can be processed in parallel with the frozen pretrained experts, capitalizing on the inherent sparsity of the MoE architecture.
@@ C.1 Evaluation Across Different Model Architectures
[…]
C.2 Analysis on the Number of Personalized Experts
In the FLEx framework, each client finetunes a small subset of personalized experts while keeping the original, pre-trained experts active. The output of the new personalized expert is integrated with that of the original experts to generate the final output, as detailed in Equation (9).
To determine the optimal number of experts to personalize, we conducted an experiment varying this number from one to three. The results, summarized in Table 7, show that pruning additional experts yields only marginal performance gains while incurring an exponential increase in computational cost.
Table 7: Performance and computational cost comparison for finetuning a varying number of experts per layer. The experiment is conducted on the Databricks-dolly-15k dataset under the pathological non-IID setting. While performance gains are marginal, the required finetuning time increases exponentially.
Method CLF CQA IE Summ Avg Pruning Time (s)
FLEx (1 Expert) 52.10 38.23 40.13 42.09 43.13 58
FLEx (2 Experts) 51.98 38.07 41.26 42.16 43.36 3,246
FLEx (3 Experts) 52.24 37.83 41.12 42.81 43.50 87,091
By personalizing a single expert, FLEx effectively captures client-specific data patterns without imposing prohibitive computational demands. This makes our method practical for real-world federated learning scenarios, where clients typically operate with limited computational resources.
C.3 Discussion on the Gating Activation Function
Our empirical results in Section 5.3 highlight the superior performance of the sigmoid activation for the personalized expert’s gating mechanism. We attribute this to the inherent properties of the sigmoid function. Its output, bounded within the $(0,1)$ range, naturally serves as a well-calibrated, probabilistic soft switch. This allows the model to learn a smooth and interpretable weighting for engaging the personalized expert based on the input token.
In contrast, the alternative functions lack this intuitive characteristic. Specifically, the ReLU-based weighting for side expert $g_{e,j}^{l}$ can be expressed as:
$g_{e,j}^{l}$ $=\text{ReLU}(\text{router}_{e,l}(u_{j}^{l})).$ (10)
$g_{e,j}^{l}$ $=\text{Tanh}(\text{router}_{e,l}(u_{j}^{l})).$ (11)
The unbounded nature of the ReLU function’s output $([0,\infty))$ can introduce training instability when used as a gate, as it does not represent a normalized weight. Similarly, the Tanh function’s output in the (1,1)(-1,1) range is less suitable for a gating mechanism, as negative values do not have a clear interpretation for modulating an expert’s contribution. Therefore, the sigmoid function provides an inductive bias that is better aligned with the task of dynamically blending a specialized expert with a set of generalist pretrained experts.
C.4 Comparison with Traditional Federated Learning
@@
C.5 Robustness across Diverse Knowledge Domains
To investigate whether FLEx’s advantages extend from task-based heterogeneity to domain-based heterogeneity, we conducted experiments on three domain-specific datasets: Alpaca-gpt4 (general), Finance-Alpaca (finance), and MedAlpaca (medical). The results in Table 8 further validate our initial findings. FLEx consistently outperforms all baselines across every domain, achieving the highest average ROUGE-L score. This demonstrates that our method is not only effective at managing different instruction-following tasks but is also robust in federated environments where clients represent distinct, specialized knowledge domains.
Table 8: ROUGE-L performance on domain-specific datasets (Alpaca-gpt4, Finance-Alpaca, and MedAlpaca) under a pathological non-IID setting. Best and second-best results among FL methods are bolded and underlined.
Method Alpaca-gpt4 Finance-Alpaca MedAlpaca Avg
Local Training 31.31 28.64 30.93 30.29
MoE + FedAvg 30.11 28.17 30.06 29.44
MoE + FedAvgM 28.90 27.18 28.78 28.28
MoE + FedAdam 29.23 28.70 31.18 27.70
MoE + FedAdagrad 30.24 28.79 30.60 29.87
MoE + SCAFFOLD 29.96 28.38 28.92 29.08
MoE + FedProx 30.05 28.60 30.01 29.55
MoE + FedYogi 29.73 28.42 30.51 29.55
[…]
C.6 Experiments on Expert Load Balancing
We conducted experiments to evaluate the activation frequency of experts within the Qwen1.5-MoE-A2.7B model using the C4 dataset [^4] https://huggingface.co/datasets/allenai/c4 across different methods. Figure 5 illustrates the activation count for individual experts. Specifically, the original model and the FedAvg method tend to produce uneven workloads, causing certain experts (e.g., expert 31 and 43), to become significantly overloaded. Specifically, the original model exhibits a standard deviation of 1859.80, while FedAvg demonstrates a slightly higher imbalance with a standard deviation of 1906.85. In contrast, our proposed FLEx method significantly improves load balancing among experts, achieving a notably lower standard deviation of 1259.88. This analysis demonstrates that our method achieves a more balanced workload distribution across different experts.
@@ C.6 Experiments on Expert Load Balancing
C.7 Data Distribution for Non-IID
@@
Figure 6: Distribution of instruction types across different clients in federated learning under various settings. The first panel shows the pathological distribution, where each client has a highly imbalanced data distribution. The second panel illustrates the Dirichlet distribution with $\alpha$=0.1, resulting in a more skewed distribution across clients. The third panel presents the Dirichlet distribution with $\alpha$=1.0, demonstrating a more uniform distribution of categories (classification, closed question answering, information extraction, and summarization) across clients.
C.8 Overview of Hyperparameters
@@
C.9 The ROUGE-L Metric
ROUGE-L (Recall-Oriented Understudy for Gisting Evaluation - Longest Common Subsequence) is a metric that evaluates a candidate text by measuring its longest common subsequence (LCS). An LCS is the longest sequence of words that is shared between two texts, maintaining their relative order but not requiring that the words be contiguous.
The metric is calculated using LCS-based recall ($R_{lcs}$), precision ($P_{lcs}$), and their F-score ($F_{lcs}$). Given a reference text $X$ of length $m$ and a candidate text $Y$ of length $n$, these components are defined as:
$R_{lcs}=\frac{LCS(X,Y)}{m},\qquad P_{lcs}=\frac{LCS(X,Y)}{n},$
where $LCS(X,Y)$ denotes the length of the longest common subsequence. The final ROUGE-L score is the F-score, which is the harmonic mean of precision and recall:
$F_{lcs}=\frac{(1+\beta^{2})R_{lcs}P_{lcs}}{R_{lcs}+\beta^{2}P_{lcs}}.$
The parameter $\beta$ controls the relative importance of recall versus precision. When $\beta=1$, it becomes the standard F1-score, giving equal weight to both.
C.10 Dataset Information
@@ D.3 Vicuna Harmless Template
Appendix E Use of LLMs
A large language model was used to assist with the language refinement of this manuscript. Specifically, the LLM was employed to improve the grammar, clarity, and style of the text in the Methodology and Experiments sections. The core scientific ideas, experimental design, and interpretation of results were conceived and articulated entirely by the human authors. The authors have thoroughly reviewed and edited all AI-generated suggestions and assume full responsibility for the final content and its scientific integrity.